OpenStudy (amonoconnor):

I quite honestly am not sure how to start this one... Find y' if 3 = (cos(x))^sin(y). Any and all help is greatly appreciated!

OpenStudy (freckles):

take ln( ) of both sides

OpenStudy (freckles):

this will allow you to bring that sin(y) down by using the power rule for log

OpenStudy (amonoconnor):

Is this an Implicit Differentiation situation?

OpenStudy (freckles):

it is an implicit equation so yes

OpenStudy (amonoconnor):

Thank you! Alright... I'm gonna dive in, and see what I can get as an answer!

OpenStudy (amonoconnor):

Okay, so I got: (and a snack) \[\frac{1}{3} = \sin(y)*(-\tan(x))\]

OpenStudy (amonoconnor):

So... what's "y prime" if it doesn't start off equal to y?

OpenStudy (alekos):

No, I don't think that's quite right

OpenStudy (freckles):

@amonoconnor do you know chain rule?

OpenStudy (freckles):

also ln(3) doesn't equal 1/3 and it looks like you completely missed that there was a product on the right hand side

OpenStudy (freckles):

oh but it being ln(3) or 1/3 won't matter in the end because the derivative of ln(3) or even 1/3 is 0 but yeah it should be ln(3) on the left hand side before differentiating and then 0 after differentiating

OpenStudy (amonoconnor):

Is this right? \[3 = (\cos(x))^{\sin(y)}\] \[\ln3 = \ln[\cos(x)^{\sin(y)}]\] \[\ln3 = \sin(y) * \ln(\cos(x))\] Now taking the derivative... \[(d/dx(\ln3) = 1/x = 1/3) = (\cos(y)) + (\frac{1}{\cos(x)})*(-\sin(x))\] \[\frac{1}{3} = \cos(y) + (-\tan(x))\] \[1/3 = \cos(y) - \tan(x)\]

OpenStudy (freckles):

but d/dx(ln(3)) is 0 you cannot treat 3 like it is x ...

OpenStudy (amonoconnor):

so... \[d/dx(lnx) = \frac{1}{x}\] does't apply?

OpenStudy (freckles):

also your chain rule part is missing when you differentiated sin(y) you also did not bring down the parts of the product rule you are for some reason saying (fg)'=f'+g' while this is not true (fg)'=f'g+g'f

OpenStudy (freckles):

3 is not x 3 is a constant

OpenStudy (freckles):

d/dx (constant)=0 d/dx( ln(3))=0 since ln(3) is constant

OpenStudy (amonoconnor):

Gotcha... okay, let me try again. :)

OpenStudy (freckles):

\[\ln(3)=\sin(y) \cdot \ln(\cos(x)) \\ \frac{d}{dx}(\ln(3))=\frac{d}{dx}[ \sin(y) \cdot \ln(\cos(x)] \] try using constant rule on left and product rule on right hand side and for both factors in the product you will need to use chain rule you are differentiating ln(cos(x)) correctly but not sin(y)

OpenStudy (freckles):

\[\frac{d}{dx} \sin(y)=y' \cos(y)\]

OpenStudy (amonoconnor):

\[y' = \frac{\sin(y)\tan(x)}{\cos(y)*\ln(\cos(x))}\]

OpenStudy (freckles):

\[0=\ln(\cos(x) ) \cdot \frac{d}{dx} (\sin(y))+ \sin(y) \cdot \frac{d}{dx}(\ln(\cos(x)) \\ 0 =\ln(\cos(x)) \cdot y' \cos(y)+\sin(y) \cdot \frac{-\sin(x)}{\cos(x)}\] and then solve for y' this will involving adding something on both sides then dividing something on both sides oh wow nevermind looks likes you did it

OpenStudy (freckles):

looks awesome

OpenStudy (amonoconnor):

Thank You!!!

OpenStudy (freckles):

you could do one more thing

OpenStudy (freckles):

but you are answer is right

OpenStudy (amonoconnor):

What else!? :)

OpenStudy (freckles):

\[y'=\frac{\sin(y) \tan(x)}{\cos(y) \cdot \ln(\cos(x))}= \frac{\frac{\sin(y)}{\cos(y)} \tan(x)}{\ln(\cos(x))} =\frac{\tan(y) \tan(x)}{\ln(\cos(x))}\]

OpenStudy (freckles):

but you don't have to do that

OpenStudy (freckles):

do you have any questions on this question/

OpenStudy (amonoconnor):

Nope, not anymore:) Thank you again!

OpenStudy (freckles):

kk cool stuff

OpenStudy (alekos):

well done connor

OpenStudy (amonoconnor):

Thank you! :)